OpenStudy (anonymous):

Given that f(x) = x^{7} h(x) h(-1) = 5 h'(-1) = 8 What is f'( -1 )? How would i find f'(-1)

OpenStudy (luigi0210):

f'(x) as derivative right?

OpenStudy (anonymous):

\[f(x)=x^7h(x)~~\Rightarrow~~f'(x)=7x^6h(x)+x^7h'(x)\] by the product rule. Given that \(h(-1)=5\) and \(h'(-1)=8\), you have \[f'(-1)=7(-1)^6h(-1)+(-1)^7h'(-1)\\ f'(-1)=-7(1)(5)+(-1)(8)\]